a rectangular field is 85 long and 50 meters wide. give the length and width of another rectangular field that has the same perimeter but a larger area...

Answer by checkley79(3341)

AREA=85*50=4,250 M^2
P=2L+2W
P=2*85+2*50=170+100=270 M. FOR THE PERIMETER.
A SQUARE FIELD OD SIDES=67.5 IS LARGER.
AREA=67.5^2=4,556.25 M^2 FOE THE SQUARE.
P=67.5*4=270 FOR THE NEW PERIMETER.